# wav2gloss: Generating Interlinear Glossed Text from Speech

### v1

@@ Abstract
Thousands of the world’s languages are in danger of extinction—a tremendous threat to cultural identities and human language diversity. Interlinear Glossed Text (IGT) is a form of linguistic annotation that can support documentation and resource creation for these languages’ communities. IGT typically consists of (1) transcriptions, (2) morphological segmentation, (3) glosses, and (4) free translations to a majority language. We propose wav2gloss: a task to extract these four annotation components automatically from speech, and introduce the first dataset to this end, Fieldwork: [^1] https://huggingface.co/datasets/wav2gloss/fieldwork a corpus of speech with all these annotations covering 37 languages with standard formatting and train/dev/test splits. We compare end-to-end and cascaded wav2gloss methods, with analysis suggesting that pre-trained decoders assist with translation and glossing, that multi-task and multilingual approaches are underperformant, and that end-to-end systems perform better than cascaded systems, despite the text-only systems’ advantages. We provide benchmarks to lay the ground work for future research on IGT generation from speech.
@@ 1 Introduction
This is a tractable problem. While linguists can do little or nothing to address the underlying causes of language endangerment—which form an intersecting lattice of political, cultural, and economic factors—speech and natural language processing researchers can do even less. Technologists can, however, facilitate the efforts of field linguists and language workers to document endangered languages by developing technologies that make the mammoth tasks of annotating field data surmountable (Shi et al., 2021a, b). For example, they can develop models that automatically generate first-pass transcriptions from raw speech. This practice has demonstrated dramatic speed-ups for human annotators Amith et al. (2021).
In direct support of this goal, we propose a new speech and language processing task: wav2gloss. This task assumes recorded speech as the only input. The output consists of aligned annotations for transcription (with and without segmentation), glossing, and translation. In order to allow the research community to participate in this task, we introduce the following:
@@
    Benchmarks based on well known speech and NLP models, including end-to-end and cascaded approaches to predict IGT from speech.
@@ 2 Dataset
Glottocode Name $C$ Type Train (h) Dev+Test (h)
@@
Our first step in adapting the Fieldwork corpora for our four subtasks is to select languages where both audio and IGT are available. There are some overlaps between DoReCo, Multi-CAST, and INEL. For each language that appears more than once in those three sources, we only use the data from the source that contains the highest number of utterances. We do not attempt to merge corpora of the same language from multiple sources. We also require all languages we select to have a permissive $C$ license without a No Derivatives (ND) restriction. [^4] We base our decision on a layperson’s reading of the Creative Commons licenses, which count cleaned, reformatted, and standardized versions of datasets as derivative works. The list of language corpora selected along with their license information, and the number of hours of data available in training and combined dev and test splits is shown in Table 1. See Figure 2 for an overview of the data processing pipelines.
Annotated data come in a variety of formats such as JSON or XML. Most of our source data come in XML-based formats—with ELAN (Brugman and Russel, 2004) being the most popular—as well as EXMARaLDA (Schmidt and Wörner, 2014) and Pangloss DTD (Michailovsky and Jacobson, 2001). [^5] See von Prince and Nordhoff (2020) for a more detailed overview ELAN and common annotator practices. In ELAN and EXMARaLDA, annotations are text strings with beginning and end time stamps, organized into tiers including underlying form, surface form, transcription, gloss, and unique ID.
@@
Documentation Reference Corpus (Paschen et al., 2020) provides time-aligned transcriptions for 51 under-resourced languages. Paschen et al. processed each language corpus through a pipeline with consistency checking, multiple rounds of audio-text alignment, and manual corrections. We selected only data that is marked as fully (vs. partially or not) glossed. The annotations in the DoReCo dataset are in the ELAN format and organized such that transcriptions, translations, and utterance ID share the same time span at the utterance level, and the underlying forms and glosses share the time span at the morpheme level within utterance time spans. After extracting the annotations, we drop words or utterances that consist only of the pause marker <p:>, and replace each span marked as <label<text>> with its text component.
@@
Grammars, Corpora and Language Technology for Indigenous Northern Eurasian Languages is an ongoing project at the Academy of Sciences and Humanities in Hamburg and the University of Hamburg that focuses on gathering resources for indigenous languages and language varieties of Northern Eurasia. [^8] https://www.slm.uni-hamburg.de/en/inel.html We use all four of the languages released so far (see Table 1) and use all annotated speech available. The datasets are annotated with EXMARaLDA, which is structurally similar to ELAN.
@@
Ainu is a nearly extinct language spoken in Hokkaido, Japan. The NINJAL Glossed Audio Corpus of Ainu Folklore (Nakagawa et al., 2021) contains recordings of 38 traditional Ainu folktales by two Ainu speakers, along with their transcriptions (in Latin script and katakana), English and Japanese translations, and underlying and surface gloss forms in English and Japanese. We used the Latin transcriptions and English translation/glosses. We scraped data via the corpus’s web interface [^11] https://ainu.ninjal.ac.jp/folklore/en/ and stored them in their native JSON format. We communicated with the authors and obtained permission to share them.
@@ 3 Experiments
We provide benchmarks for automatically generating IGT by fine-tuning pre-trained speech and text models. We choose commonly used models and methods with their corresponding codebases to provide a solid baseline for future research. See Appendix A for details on models sizes and hyper-parameter settings.
@@ 3.1 End-to-End models
For end-to-end approaches, we use ESPnet Watanabe et al. (2018) to employ two representative families of state-of-the-art pre-trained speech models: self-supervised and semi-supervised. The first self-supervised model we employ is WavLM Large Chen et al. (2022), which achieves state-of-the-art performance on the SUPERB benchmark, a leaderboard for various speech-related tasks Yang et al. (2021); Feng et al. (2023). We also use XLS-R-300M Babu et al. (2021), a model specifically trained for cross-lingual capabilities, which has shown superior performance on the ML-SUPERB benchmark in multilingual tasks Shi et al. (2023a, b). WavLM and XLS-R are also in the family of HuBERT- and wav2vec2.0-like models Hsu et al. (2021); Baevski et al. (2020), respectively, which are commonly used self-supervised models. Similar to Chen et al. (2023), while we freeze the self-supervised models to preserve acoustic knowledge, we attach a conformer encoder and transformer decoder Guo et al. (2021) to support the four different annotations we infer. Both the encoder (50M parameters) and decoder (26M parameters) have six blocks and eight attention heads, with the addition of the 315M frozen parameters from each of the pre-trained models, bringing the total parameter count to 391M. We train the model with CTC-Attention loss Kim et al. (2017), where CTC and attention loss are applied to the encoder and decoder, respectively. We train the language model from scratch and employ character-level tokenization with added language and task tokens. For details, see our public source code. [^12] https://anonymous.4open.science/r/espnet-4B6F/egs2/wav2gloss/asr1/
On the other hand, supervised speech models, such as Whisper Radford et al. (2023) or OWSM Peng et al. (2023), show reasonably robust performance in various tasks, especially ASR. Hence, we fine-tune the OWSM-v3.1-base model (Peng et al. 2024, 101M parameters), which is an open-sourced reproduction of Whisper with public training software and datasets. We use OWSM in our experiments because its open-source nature is desireable for reproducibility and thorough scientific analysis. These supervised models already contain a predefined vocabulary and corresponding tokenizer. OWSM uses a BPE tokenizer with 50k vocabulary, trained on a random subset of its training data. We add a specific token per language and two task tokens for gloss and underlying forms. For transcription and translation, we utilize the existing task tokens. Just as with self-supervised models, we fine-tune with the CTC-Attention loss (the same approach as in OWSM pre-training). The source code for fine-tuning OWSM is available in a public repository. [^13] https://anonymous.4open.science/r/finetune_owsm-F8AA/
@@ 3.2 Cascaded model
From our preliminary experiments, we see that predicting glosses from speech is much more challenging than predicting transcriptions (i.e. ASR). The 2023 SIGMORPHON shared task on interlinear glossing (Ginn et al., 2023) showed that text-to-gloss models can achieve reasonable gloss prediction accuracy. Therefore, we evaluate a cascaded approach where we take the best performing ASR models in the end-to-end setting, and use their transcription outputs as inputs to a text-to-gloss model. We use two text models initialized from ByT5-base (Xue et al. 2022, 582M parameters), one trained only on Fieldwork for the underlying form, gloss, and translation tasks; the other fine-tuned first on ODIN (Lewis and Xia, 2010) and then fine-tuned on Fieldwork for gloss and translation, similar to the approach of He et al. (2023). While the ByT5 model was not the best performing in the shared task, it can easily support multilingual training and does not strictly enforce an alignment between input and output tokens, which theoretically makes it less sensitive to noise, and more robust for a cascaded approach. Preliminary experiments have also shown that ByT5 is better at handling unsegmented inputs than other approaches.
@@
Table 2: Results from multilingual experiments. The languages are split into two groups: “seen”, where the languages are in the training set, and “unseen”, where the languages are not in the training set. Each number represent an average of that metric across the languages in that group (macro-averaging).
@@ 5 Discussion
Some trends in our experimental results highlight trends that may assist further development of wav2gloss technologies.
@@
Our cascaded models do not fully avail text-based potential
From the results of the models trained on ground truth text in Table 2, it is clear that annotation inference is easier from text than from speech. However, the advantage is not enough to overcome error propagation introduced by a cascade approach, at least for glossing. One appeal of text-input models is training data availability, since text IGT data is far more plentiful than speech. Our own results—indicating that fine-tuning on noisy IGT (ODIN) improves glossing and translation—demonstrate some of this potential. Novel approaches to this task should take this into account, and perhaps employ a multimodal model accepting both speech and text inputs.
Figure 3: Comparing the seen language performance of multilingual transcription OWSM E2E model with monolingual models, trained separately with individual languages. Multilingual and monolingual performance is denoted as dark blue and light pink dots, respectively. Same language is connected with the black line.
@@
Multilingual training can vitally boost model performance for low-resource languages in some settings (Chen et al., 2019). However, multilinguality can also cause performance degradation for some languages, a phenomenon known as the “curse of multilinguality” Conneau et al. (2020). Given this theoretical tention, we analyze monolingual versus multilingual training directly. We can only fairly compare OWSM models in this analysis, since self-supervised models rely on language models trained from scratch, and hence have different vocabulary sizes in monolingual vs. multilingual settings. Transcription error rates from end-to-end single-task OWSM models using monolingual versus multilingual training are shown in Figure 3. We observe that CER improves (decreases) as the dataset size increases. Generally, multilingual models perform better for small datasets, but this trend inverts as data increase, indicating multilingual performance degradation. This performance degradation is less pronounced for the highest-resource languages (Dolgan, Kamas, and Ainu; on the far right in Figure 3) than for other languages. We suspect that multilinguality is less harmful for these languages due to our checkpoint saving strategy. (See Section 3.1.) Because we keep the checkpoint with the best dev set accuracy in multilingual training, the checkpoints are biased towards languages with the highest dev set representation.
@@ 7 Conclusion and Future Work
In this work, we make a number of strides to advance technology-assisted documentation needed by language communities. We define a new task, wav2gloss, to predict IGT annotations directly from speech. We present Fieldwork, the first dataset for this task, comprised of audio files and expert annotations that were cleaned and formatted. We provide benchmarks across various training methodologies for transcription, underlying form prediction, glossing, and translation. And we analyse various prominent trends from experimental results. These data, benchmarks, and preliminary experimental insights provide a strong foundation for future wav2gloss breakthroughs, to expedite creation of needed language resources for communities of dying languages.
This work may be continued in a number of ways. Future researchers may further normalize IGT labels, to expand Fieldwork to more diverse languages and phenomena. This is labor intensive, as previous research (List et al., 2021) and we ourselves observed, given inconsistent use of labels and language-specific phenomena across IGT collections. Expansion of Fieldwork to more languages may also come through community-driven projects to benefit low-resource language communities and academics. Researchers may also focus on improving our work’s modeling capabilities—e.g., adapting models to zero-shot performance. In our own work, all models perform notably better on seen than on unseen languages. Future work may mitigate this by mapping all transcriptions to a shared vocabulary, such as IPA, to minimize superficial language differences.
@@
While we expect our contributions to be of significant value to the research community, we wish to acknowledge a significant limitations to consider. Most notably to start, as is apparent from the scores in Table 2, our models’ performance in any of the four subtasks is not sufficient to render useful outputs for application. This highlights the challenging nature of working with low-resource languages and the novel wav2gloss task. By the release of our dataset and benchmarks, we mean to spur future iterations that improve upon our results and move towards applicable solutions for low-resource language communities.
We acknowledge also that, since we did not perform linguistic annotations ourselves of the IGT datasets we include, our dataset’s quality is tied to the accuracy of others’ linguistic expertise and annotations. Since errors in annotation for training data, such as misalignments and mislabels, can cause mild to severe errors in machine learning outputs, we encourage users of our data to proceed with caution. We ensure our data’s quality via cleaning and filtering, in addition to random manual inspection finding no severe errors, but we encouraged continued vigilance in this vein.

### v2

@@ Abstract
Thousands of the world’s languages are in danger of extinction—a tremendous threat to cultural identities and human language diversity. Interlinear Glossed Text (IGT) is a form of linguistic annotation that can support documentation and resource creation for these languages’ communities. IGT typically consists of (1) transcriptions, (2) morphological segmentation, (3) glosses, and (4) free translations to a majority language. We propose wav2gloss: a task in which these four annotation components are extracted automatically from speech, and introduce the first dataset to this end, Fieldwork: [^1] https://huggingface.co/datasets/wav2gloss/fieldwork a corpus of speech with all these annotations, derived from the work of field linguists, covering 37 languages, with standard formatting, and train/dev/test splits. We provide various baselines to lay the groundwork for future research on IGT generation from speech, such as end-to-end versus cascaded, monolingual versus multilingual, and single-task versus multi-task approaches.
@@ 1 Introduction
Producing IGT from audio is a tractable problem. While linguists can do little or nothing to address the underlying causes of language endangerment—which form an intersecting lattice of political, cultural, and economic factors—speech and natural language processing researchers can do even less. Technologists can, however, facilitate the efforts of field linguists and language workers to document endangered languages by developing technologies that make the mammoth tasks of annotating field data surmountable (Shi et al., 2021a, b). For example, they can develop models that automatically generate first-pass transcriptions from raw speech. Research has demonstrated that such models can speed up transcription dramatically Amith et al. (2021).
In direct support of the goal of language documentation, we propose a new speech and language processing task: wav2gloss. This task assumes recorded speech as the only input. The output consists of aligned annotations for transcription (with and without segmentation), glossing, and translation. In order to allow the research community to participate in this task, we introduce the following:
@@
    Benchmarks based on well known speech and NLP models, including end-to-end and cascaded approaches to predicting IGT from speech.
@@ 2 Dataset
Glottocode Name CC Type Train (h) Dev+Test (h)
@@
Our first step in adapting the Fieldwork corpora for our four subtasks is to select languages where both audio and IGT are available. There are some overlaps between DoReCo, Multi-CAST, and INEL. For each language that appears more than once in those three sources, we only use the data from the source that contains the highest number of utterances. We do not attempt to merge corpora of the same language from multiple sources. We also require all languages we select to have a permissive CC license without a No Derivatives (ND) restriction. [^4] We base our decision on a layperson’s reading of the Creative Commons licenses, which counts cleaned, reformatted, and standardized versions of datasets as derivative works. The list of language corpora selected along with their license information, and the number of hours of data available in training and combined dev and test splits is shown in Table 1. See Figure 2 for an overview of the data processing pipelines.
Annotated data come in a variety of formats such as JSON or XML. Most of our source data come in XML-based formats—with ELAN (Brugman and Russel, 2004) being the most popular—as well as EXMARaLDA (Schmidt and Wörner, 2014) and Pangloss DTD (Michailovsky and Jacobson, 2001). [^5] See von Prince and Nordhoff (2020) for a more detailed overview of ELAN and common annotator practices. In ELAN and EXMARaLDA, annotations are text strings with beginning and end time stamps, organized into tiers including underlying form, surface form, transcription, gloss, and unique ID.
@@
Documentation Reference Corpus (Paschen et al., 2020) provides time-aligned transcriptions for 51 under-resourced languages. Paschen et al. processed each language corpus through a pipeline with consistency checking, multiple rounds of audio-text alignment, and manual corrections. We selected only data that is marked as fully (vs. partially or not) glossed. The annotations in the DoReCo dataset are in the ELAN format and organized such that transcriptions, translations, and utterance IDs share the same time span at the utterance level, and the underlying forms and glosses share the time span at the morpheme level within utterance time spans. After extracting the annotations, we drop words or utterances that consist only of the pause marker <p:>, and replace each span marked as <label<text>> with its text component.
@@
Grammars, Corpora and Language Technology for Indigenous Northern Eurasian Languages is an ongoing project at the Academy of Sciences and Humanities in Hamburg and the University of Hamburg that focuses on gathering resources for indigenous languages and language varieties of Northern Eurasia. [^8] https://www.slm.uni-hamburg.de/en/inel.html We use all four of the languages released so far (see Table 1) and use all annotated speech available. The datasets are annotated with EXMARaLDA (Schmid and Wörner, 2014), which is structurally similar to ELAN.
@@
Ainu is a nearly extinct language spoken in Hokkaido, Japan. The NINJAL Glossed Audio Corpus of Ainu Folklore (Nakagawa et al., 2021) contains recordings of 38 traditional Ainu folktales by two Ainu speakers, along with their transcriptions (in Latin script and occasionally Japanese script), English and Japanese translations, and underlying and surface gloss forms in English and Japanese. We used the Latin transcriptions and English translation/glosses. We scraped data via the corpus’s web interface [^11] https://ainu.ninjal.ac.jp/folklore/en/ and stored them in their native JSON format. We communicated with the authors and obtained permission to share them.
@@ 3 Experiments
We provide benchmarks for automatically generating IGT by fine-tuning pre-trained speech and text models. We choose commonly used models, methods, and finetuning settings with their corresponding codebases to provide a solid baseline for future research. See Appendix A for details on models sizes and hyper-parameter settings.
@@ 3.1 End-to-End models
For end-to-end approaches, we use ESPnet Watanabe et al. (2018) to employ two representative families of state-of-the-art pre-trained speech models: self-supervised and semi-supervised. The first self-supervised model we employ is WavLM Large Chen et al. (2022), which achieves state-of-the-art performance on the SUPERB benchmark, a leaderboard for various speech-related tasks Yang et al. (2021); Feng et al. (2023). We also use XLS-R-300M Babu et al. (2021), a model specifically trained for cross-lingual capabilities, which has shown superior performance on the ML-SUPERB benchmark in multilingual tasks Shi et al. (2023a, b). WavLM and XLS-R are also in the family of HuBERT- and wav2vec2.0-like models Hsu et al. (2021); Baevski et al. (2020), respectively, which are commonly used self-supervised models. Similar to Chen et al. (2023), while we freeze the self-supervised models to preserve acoustic knowledge, we attach a conformer encoder and transformer decoder Guo et al. (2021) to support the four different annotations we infer. Both the encoder (50M parameters) and decoder (26M parameters) have six blocks and eight attention heads, with the addition of the 315M frozen parameters from each of the pre-trained models, bringing the total parameter count to 391M. WavLM was pretrained on English only, while XLS-R was pretrained on multiple languages, two of which—Mandarin and Persian—are also in Fieldwork. We train the model with CTC-Attention loss Kim et al. (2017), where CTC and attention loss are applied to the encoder and decoder, respectively. We train the language model from scratch and employ character-level tokenization with added language and task tokens. For details, see our public source code. [^12] https://github.com/juice500ml/espnet/tree/wav2gloss/egs2/wav2gloss/asr1
On the other hand, supervised speech models, such as Whisper Radford et al. (2023) or OWSM Peng et al. (2023), show reasonably robust performance in various tasks, especially ASR. Hence, we fine-tune the OWSM-v3.1-base model (Peng et al. 2024, 101M parameters), which is an open-sourced reproduction of Whisper with public training software and datasets. We use OWSM in our experiments because its open-source nature is desirable for reproducibility and thorough scientific analysis. These supervised models already contain a predefined vocabulary and the corresponding tokenizer. OWSM uses a BPE tokenizer with 50k vocabulary, trained on a random subset of its training data. We add a specific token per language and two task tokens for gloss and underlying forms. For transcription and translation, we utilize the existing task tokens. The pretraining corpus of OWSM is also multilingual, with Mandarin and Persian being the only two overlapping languages with Fieldwork. Just as with self-supervised models, we fine-tune with the CTC-Attention loss (the same approach as in OWSM pre-training). Similar to Rouditchenko et al. (2023), we fully fine-tune the OWSM model. The source code for fine-tuning OWSM is available in a public repository. [^13] https://github.com/juice500ml/finetune_owsm
@@ 3.2 Cascaded model
From our preliminary experiments, we see that predicting glosses from speech is much more challenging than predicting transcriptions (i.e. ASR). The 2023 SIGMORPHON shared task on interlinear glossing (Ginn et al., 2023) showed that text-to-gloss models can achieve over 90% gloss prediction accuracy in certain languages, given enough training data. Therefore, we evaluate a cascaded approach where we take the best performing ASR models in the end-to-end setting, and use their transcription outputs as inputs to a text-to-gloss model. We use two text models initialized from ByT5-base (Xue et al. 2022, 582M parameters), one trained only on Fieldwork for the underlying form, gloss, and translation tasks; the other fine-tuned first on ODIN (Lewis and Xia, 2010) and then fine-tuned on Fieldwork for gloss and translation, similar to the approach of He et al. (2023). All text models are single-task, meaning we train a separate model for each task within each setting. While the ByT5 model was not the best performing in the shared task, it can easily support multilingual training and does not require the inputs to be segmented, and therefore is more suitable for the unsegmented outputs of the ASR systems.
@@
Table 2: Results from multilingual experiments. The languages are split into two groups: “seen”, where the languages are in the training set, and “unseen”, where the languages are not in the training set. Each number represents an average of that metric across the languages in that group (macro-averaging). WavLM and XLS-R models have pretrained encoders while OWSM have pretrained encoder and decoder.
@@ 5 Discussion
Some aspects of our experimental results highlight trends that may assist further development of wav2gloss technologies.
@@
Our cascaded models do not fully realize text-based potential
From the results of the models trained on ground truth text in Table 2, it is clear that annotation inference is easier from text than from speech. However, the advantage is not enough to overcome error propagation introduced by a cascade approach, at least for glossing. One appeal of text-input models is training data availability, since text IGT data is far more plentiful than speech. Our own results—indicating that fine-tuning on noisy IGT (ODIN) improves glossing and translation—demonstrate some of this potential. For machine translation in particular, there are specialized pretrained models with much higher multilingual MT performance that could be deployed as part of the pipeline. Novel approaches to this task should take this into account, and perhaps employ a multimodal model accepting both speech and text inputs.
Figure 3: Comparing the seen language performance of multilingual transcription OWSM E2E model with monolingual models, trained separately with individual languages. Multilingual and monolingual performance is denoted as dark blue and light pink dots, respectively. Same language is connected with a black line.
@@
Multilingual training can vitally boost model performance for low-resource languages in some settings (Chen et al., 2019). However, multilinguality can also cause performance degradation for some languages, a phenomenon known as the “curse of multilinguality” Conneau et al. (2020). Given this theoretical tension, we analyze monolingual versus multilingual training directly. We can only fairly compare OWSM models in this analysis, since self-supervised models rely on language models trained from scratch, and hence have different vocabulary sizes in monolingual vs. multilingual settings. Transcription error rates from end-to-end single-task OWSM models using monolingual versus multilingual training are shown in Figure 3. We observe that CER improves (decreases) as the dataset size increases. Generally, multilingual models perform better for small datasets, but this trend inverts as data increase, indicating multilingual performance degradation. This performance degradation is less pronounced for the highest-resource languages (Dolgan, Kamas, and Ainu; on the far right in Figure 3) than for other languages. We suspect that multilinguality is less harmful for these languages due to our checkpoint saving strategy. (See Section 3.1.) Because we keep the checkpoint with the best dev set accuracy in multilingual training, the checkpoints are biased towards languages with the highest dev set representation.
@@ 7 Conclusion and Future Work
In this work, we make a number of strides to advance technology-assisted documentation needed by language communities. We define a new task, wav2gloss, which is to predict IGT annotations directly from speech. We present Fieldwork, the first dataset for this task, comprised of audio files and expert annotations that were cleaned and formatted. We provide benchmarks across various training methodologies for transcription, underlying form prediction, glossing, and translation. And we analyze various prominent trends from experimental results. These data, benchmarks, and preliminary experimental insights provide a strong foundation for future wav2gloss breakthroughs, to expedite creation of needed language resources for communities of dying languages.
This work may be continued in a number of ways. The models presented in this work represent baseline approaches, and we hope this work will spur on more experimentation on the optimal hyperparameters for this task. In particular, the choice of model size and the tokenization method can be further explored. Moreover, we think multi-modal setups such as Barrault et al. (2023) where both speech and text inputs can be fed into the model at the same time can be very promising for wav2gloss. Another possible avenue of research is the use of cascades consisting of more than two models, for example, ASR into morpheme segmentation into glossing.
Future researchers may also further normalize IGT labels, to expand Fieldwork to more diverse languages and phenomena. This is labor intensive, as previous research (List et al., 2021) and we ourselves observed, given inconsistent use of labels and language-specific phenomena across IGT collections. Expansion of Fieldwork to more languages may also come through community-driven projects to benefit low-resource language communities and academics. Researchers may also focus on improving our work’s modeling capabilities—e.g., adapting models to zero-shot performance. In our own work, all models perform notably better on seen than on unseen languages. Future work may mitigate this by mapping all transcriptions to a shared vocabulary, such as IPA, to minimize superficial orthographic differences.
@@
While we expect our contributions to be of significant value to the research community, we wish to acknowledge significant limitations to consider. Most notably to start, as is apparent from the scores in Table 2, our models’ performance in any of the four subtasks is not sufficient to render useful outputs for application. This highlights the challenging nature of working with low-resource languages and the novel wav2gloss task. By the release of our dataset and benchmarks, we mean to spur future iterations that improve upon our results and move towards applicable solutions for low-resource language communities.
We acknowledge also that, since we did not perform linguistic annotations ourselves of the IGT datasets we include, our dataset’s quality is tied to the accuracy of others’ linguistic expertise and annotations. Since errors in annotation for training data, such as misalignments and mislabels, can cause mild to severe errors in machine learning outputs, we encourage users of our data to proceed with caution. We ensure our data’s quality via cleaning and filtering, in addition to random manual inspection finding no severe errors, but we encourage continued vigilance in this vein.
@@
Acknowledgements
This work is supported by the US National Science Foundation grant #2211951. We also thank the linguists whose resources we used and their collaborators. Without them, this work would not have been possible. For the computational resources, we used the Bridges2 system at PSC and Delta system at NCSA through allocation CIS210014 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is also supported by the US National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. Finally, we thank Paola Garcia, the anonymous reviewers, and the Area Chair for their valuable feedback to the present paper.